### v2

@@ Abstract
Keywords:
Reinforcement Learning Navigation Unmanned Systems Contorl.
1 Introduction
Reinforcement Learning (RL) optimizes decision-making and behavior in dynamic environments through trial and error and reward mechanisms. It is widely used in fields such as gaming [1, 2, 3, 4], control [5, 6, 7], autonomous driving [8, 9], industrial automation [10, 11] and collaborative tasks [12, 13, 14]. RL involves agents interacting with the environment, learning through exploration and exploitation to achieve their goals. However, in real-world applications like the Source Term Estimation (STE) problem [15], the environment often lacks clear signals for the end of an episode or direct rewards, making it difficult for the agent to determine task completion.
@@ 1 Introduction
Statistical-based methods such as Infotaxis, Entrotaxis , and Dual Control for Exploitation and Exploration (DCEE) provide some solutions to these challenges. Infotaxis in [16, 17, 18] guides the search process by maximizing the rate of information acquisition, gradually reducing uncertainty to approximate the source of the gas. It is suitable for situations with high uncertainty regarding the gas source location but suffers from poor real-time performance and a tendency to get stuck in local optima. Entrotaxis in [19, 20] minimizes the entropy of future measurements to guide path selection, optimizing information acquisition based on real-time measurement data. It performs robustly in noisy environments but has poor adaptability in dynamic settings or sudden changes, making rapid adjustments difficult. DCEE in [21, 22] combines strategies of exploitation and exploration by dynamically adjusting the balance between the two, optimizing information acquisition and path selection. Suitable for highly variable environments, this strategy is complex to implement, requiring precise design of the trade-off mechanism between exploration and exploitation. Although these methods are effective in addressing the lack of environmental feedback and clear cessation signals, they share several common limitations: poor real-time performance, poor adaptability in rapidly changing environments, a tendency to get stuck in local optima, and complexity in implementation, particularly for the DCEE.
@@
2 Problem Formulation and Peliminaries
In a two-dimensional search area denoted as $\Omega\subseteq\mathbb{R}^{2}$, where there is an expectation of encountering a hazardous release, a robot equipped with a gas sensor is tasked with traversing the area to calculate the release parameters, also referred to as the source term $\Theta_{s}$. This data will serve as the requisite input for a convection-diffusion model [16], enabling the generation of hazard forecasts. Within this context, in an environment characterized by average wind speed $u_{s}\in\mathbb{R}^{+}$ in meters per second (m/s), wind direction $\phi_{s}$ in radians (rad), and diffusivity $d_{s}$ in meters squared per second (m2/s), we detect the concentration of a hazardous substance $x_{k}\in\mathbb{R}^{+}$ using a robot or a sensor located at position $p_{k}=[x_{k},y_{k}]^{T}\in\Omega$ in meters (m). This hazardous material stems from a source term located at $p_{s}\in\Omega$, and it is released at a rate/strength $q_{s}\in\mathbb{R}^{+}$ in grams per second (g/s) with an average lifespan of $\tau_{s}\in\mathbb{R}^{+}$ in seconds (s).
@@ 2 Problem Formulation and Peliminaries
2.1 Gaseous Diffusion Model
The average gas concentration [23], denoted as $m(p|\Theta_{s})$, for a given location of the mobile sensor at position $p_{k}$ over a time duration of $\tau_{s}$, can be computed utilizing the parameters of the source term $\Theta_{s}$ as follows:
@@ 2.1 Gaseous Diffusion Model
2.2 Sensor Models
When mobile robots equipped with sensors conduct gas concentration detection in an environment, the measurement results are influenced by both internal and external factors. Internal factors include the sensitivity and calibration issues of the sensors, as well as the robot’s internal dynamics during movement, such as vibrations and posture changes. These factors can significantly impact gas concentration measurements. For example, the sensitivity of metal–oxide gas sensors [24, 25, 26] can vary with atmospheric contaminants, causing fluctuations in resistance and, consequently, the voltage readings. Furthermore, calibration issues, as discussed in the context of uncalibrated sensors, can lead to inaccuracies in the measured gas concentrations.
@@ 2.2 Sensor Models
Specifically, the noise due to internal factors from the sensor is represented as $v_{\text{inter}}$. Therefore, the value of the gas concentration $c(p_{k}|\Theta)$ measured by the robot at position $p_{k}$ is:
$c(p_{k}|\Theta)=m(p_{k}|\Theta)+v_{\text{inter}},$
where $v_{\text{inter}}\sim\mathcal{N}(u,\sigma)$ and $v_{\text{inter}}$ follows a normal distribution $\mathcal{N}$ with mean $u$ and variance $\sigma$. This Gaussian noise model is widely used to emulate sensor measurements in simulations, accounting for both wind turbulence and sensor errors.
@@
2.3 Partially Observable Markov Decision Process
@@
3 Methodology
3.1 Bayesian Approximation for Belief Distribution
@@
Variational inference [27] and particle filter [28] are common Bayesian inference methods. The former approximates the target distribution by optimizing a differentiable lower bound but typically requires simplifying assumptions that may not suit complex environments. The latter, on the other hand, handles nonlinear and non-Gaussian distributions more effectively and adapts to dynamic conditions without needing such assumptions. Particle filter approximate the true distribution through sufficient sampling and offer flexibility in managing multimodal distributions, tracking potential source locations and release rates, and providing real-time updates. Therefore, particle filter outperform variational inference in STE problems.
@@
The sampling weights $\{w_{k}^{i}\}_{i=1:N}$ are updated through iterative sequential importance sampling [29]. At each step, a new sample of source parameter estimates $\{\Theta_{k}^{i}\}_{i=1:N}$ is obtained from the proposal distribution $q(\Theta_{k}^{i})$. Then, the unnormalized weight $\bar{w}_{k+1}^{(i)}$ corresponding to the source term vector $\Theta_{k}$ of the particle filter is updated as follows:
@@
As the number of learning iterations increases, the weights of most particles tend to zero, leading to the degeneracy problem. To address this issue, resampling is employed, and this method is based on the Markov Chain Monte Carlo move step [30].
@@
3.2 Autonomous Goal Detection and Cessation
@@
Algorithm 1 Reinforcement Learning with AGDC
$STD=\sqrt{diag(Cov(\Theta))},$ (7)
For example, suppose the covariance matrix of the particle filter belief state is:
$Cov(\Theta)=\begin{pmatrix}\sigma_{x}^{2}&\sigma_{xy}\\ \sigma_{xy}&\sigma_{y}^{2}\end{pmatrix}$
Using diag($\cdot$)⋅ ), the standard deviations for the $x$ and $y$ coordinates can be calculated as:
$STD_{x}=\sqrt{\sigma_{x}^{2}},\quad STD_{y}=\sqrt{\sigma_{y}^{2}}$
This approach allows for an intuitive assessment of the uncertainty for each parameter. In contrast, directly calculating the standard deviation of the entire covariance matrix would require complex matrix operations, making the results less intuitive to understand. Therefore, using diag($\cdot$) for standard deviation calculation is a more effective and intuitive method. The details of the training algorithm are shown in Algorithm 1.
4 Experiment
In this paper, we utilized the STE Environment (STEenv) as the experimental platform to investigate the application of AGDC in RL. The STE Environment, described in Section 2 or [31] , comprises a Gaussian diffusion model and a sensor model. Instead of providing rewards to the Agent, it transmits the current position and the test concentration at that position as observations. Any RL algorithm can be integrated with the AGDC module to address the challenges posed by the STE problem.
4.1 Baseline Algorithms and Evaluation Metrics
Infotaxis [16] aims to minimize the predicted posterior variance of the source location. It treats the search process as an information-gathering problem, where the objective is to reduce uncertainty about the source location.
Entrotaxis [19] enforces the maximum entropy sampling principle, which directs the agent to move towards positions of maximum uncertainty to gather more informative data about the environment and the source location.
DCEE [21] integrates both exploitation and exploration by incorporating the uncertainty into the control decisions, thus allowing the robot to not only move towards the estimated target but also probe the environment to reduce uncertainty.
@@ 4.1 Baseline Algorithms and Evaluation Metrics
Table 1: Parameter Distributions for the Training Scenarios
@@
4.2 Scenario Parameterization and Evaluation
@@
4.3 Particle Numbers and Cessation Thresholds

### v3

@@ Abstract
Introduction
Reinforcement Learning (RL) optimizes decision-making and behavior in dynamic environments through trial and error and reward mechanisms. It is widely used in fields such as gaming (Mnih et al. 2015; Silver et al. 2017; Wang et al. 2024; Wang, Zhang, and Zhang 2022), robotic control (Levine et al. 2016; Zhang et al. 2021; Wang et al. 2023), autonomous driving (Feng et al. 2023; Kiran et al. 2021), industrial automation (Degrave et al. 2022; Meyes et al. 2017), preference alignment (Liu et al. 2022; Bai et al. 2023, 2024) and collaborative tasks (Li et al. 2023, 2024; Xu et al. 2024). RL involves agents interacting with the environment, learning through exploration and exploitation to achieve their goals. However, in real-world applications like the Source Term Estimation (STE) problem (Steiner and Bushe 2001), the environment often lacks clear signals for the end of an episode or direct rewards, making it difficult for the agent to determine task completion.
@@ Introduction
Statistical-based methods such as Infotaxis, Entrotaxis , and Dual Control for Exploitation and Exploration (DCEE) provide some solutions to these challenges. Infotaxis in (Vergassola, Villermaux, and Shraiman 2007; Hutchinson, Liu, and Chen 2018; Loisy and Eloy 2022) guides the search process by maximizing the rate of information acquisition, gradually reducing uncertainty to approximate the source of the gas. It is suitable for situations with high uncertainty regarding the gas source location but suffers from poor real-time performance and a tendency to get stuck in local optima. Entrotaxis in (Hutchinson, Oh, and Chen 2018; Zhao et al. 2020) minimizes the entropy of future measurements to guide path selection, optimizing information acquisition based on real-time measurement data. It performs robustly in noisy environments but has poor adaptability in dynamic settings or sudden changes, making rapid adjustments difficult. DCEE in (Chen, Rhodes, and Liu 2021; Rhodes, Liu, and Chen 2022) combines strategies of exploitation and exploration by dynamically adjusting the balance between the two, optimizing information acquisition and path selection. Suitable for highly variable environments, this strategy is complex to implement, requiring precise design of the trade-off mechanism between exploration and exploitation. Although these methods are effective in addressing the lack of environmental feedback and clear cessation signals, they share several common limitations: poor real-time performance, poor adaptability in rapidly changing environments, a tendency to get stuck in local optima, and complexity in implementation, particularly for the DCEE.
@@
Problem Formulation and Peliminaries
In a two-dimensional search area denoted as $\Omega\subseteq\mathbb{R}^{2}$, where there is an expectation of encountering a hazardous release, a robot equipped with a gas sensor is tasked with traversing the area to calculate the release parameters, also referred to as the source term $\Theta_{s}$. This data will serve as the requisite input for a convection-diffusion model (Vergassola, Villermaux, and Shraiman 2007), enabling the generation of hazard forecasts. Within this context, in an environment characterized by average wind speed $u_{s}\in\mathbb{R}^{+}$ in meters per second (m/s), wind direction $\phi_{s}$ in radians (rad), and diffusivity $d_{s}$ in meters squared per second (m2/s), we detect the concentration of a hazardous substance $x_{k}\in\mathbb{R}^{+}$ using a robot or a sensor located at position $p_{k}=[x_{k},y_{k}]^{T}\in\Omega$ in meters (m). This hazardous material stems from a source term located at $p_{s}\in\Omega$, and it is released at a rate/strength $q_{s}\in\mathbb{R}^{+}$ in grams per second (g/s) with an average lifespan of $\tau_{s}\in\mathbb{R}^{+}$ in seconds (s).
@@
Gaseous Diffusion Model
The average gas concentration (Zhao et al. 2022), denoted as $m(p|\Theta_{s})$, for a given location of the mobile sensor at position $p_{k}$ over a time duration of $\tau_{s}$, can be computed utilizing the parameters of the source term $\Theta_{s}$ as follows:
@@
Sensor Models
When mobile robots equipped with sensors conduct gas concentration detection in an environment, the measurement results are influenced by both internal and external factors. Internal factors include the sensitivity and calibration issues of the sensors, as well as the robot’s internal dynamics during movement, such as vibrations and posture changes. These factors can significantly impact gas concentration measurements. For example, the sensitivity of metal–oxide gas sensors (Li et al. 2011; Neumann et al. 2013; Neumann and Bartholmai 2015) can vary with atmospheric contaminants, causing fluctuations in resistance and, consequently, the voltage readings. Furthermore, calibration issues, as discussed in the context of uncalibrated sensors, can lead to inaccuracies in the measured gas concentrations.
@@
Specifically, the noise due to internal factors from the sensor is represented as $v_{\text{inter}}$. Therefore, the value of the gas concentration $c(p_{k}|\Theta)$ measured by the robot at position $p_{k}$ is $c(p_{k}|\Theta)=m(p_{k}|\Theta)+v_{\text{inter}}$, where $v_{\text{inter}}\sim\mathcal{N}(u,\sigma)$ and $v_{\text{inter}}$ follows a normal distribution $\mathcal{N}$ with mean $u$ and variance $\sigma$. This Gaussian noise model is widely used to emulate sensor measurements in simulations, accounting for both wind turbulence and sensor errors.
@@
Partially Observable Markov Decision Process
@@
Methodology
Bayesian Approximation for Belief Distribution
@@
Variational inference (Jordan et al. 1999) and particle filter (Gordon, Salmond, and Smith 1993) are common Bayesian inference methods. The former approximates the target distribution by optimizing a differentiable lower bound but typically requires simplifying assumptions that may not suit complex environments. The latter, on the other hand, handles nonlinear and non-Gaussian distributions more effectively and adapts to dynamic conditions without needing such assumptions. Particle filter approximate the true distribution through sufficient sampling and offer flexibility in managing multimodal distributions, tracking potential source locations and release rates, and providing real-time updates. Therefore, particle filter outperform variational inference in STE problems.
@@
The sampling weights $\{w_{k}^{i}\}_{i=1:N}$ are updated through iterative sequential importance sampling (Tokdar and Kass 2010). At each step, a new sample of source parameter estimates $\{\Theta_{k}^{i}\}_{i=1:N}$ is obtained from the proposal distribution $q(\Theta_{k}^{i})$. Then, the unnormalized weight $\bar{w}_{k+1}^{(i)}$ corresponding to the source term vector $\Theta_{k}$ of the particle filter is updated as follows:
@@
As the number of learning iterations increases, the weights of most particles tend to zero, leading to the degeneracy problem. To address this issue, resampling is employed, and this method is based on the Markov Chain Monte Carlo move step (Ristic, Arulampalam, and Gordon 2004).
@@
Autonomous Goal Detection and Cessation
@@
$STD=\sqrt{diag(Cov(\Theta))},$ (7)
Algorithm 1 Reinforcement Learning with AGDC
@@
For example, suppose the covariance matrix of the particle filter belief state is $Cov(\Theta)=\begin{pmatrix}\sigma_{x}^{2}&\sigma_{xy}\\ \sigma_{xy}&\sigma_{y}^{2}\end{pmatrix}$. Using diag($\cdot$)⋅ ), the standard deviations for the $x$ and $y$ coordinates can be calculated as $STD_{x/y}=\sqrt{\sigma_{x/y}^{2}}$. This approach allows for an intuitive assessment of the uncertainty for each parameter. In contrast, directly calculating the standard deviation of the entire covariance matrix would require complex matrix operations, making the results less intuitive to understand. Therefore, using diag($\cdot$) for standard deviation calculation is a more effective and intuitive method. The details of the training algorithm are shown in Algorithm 1.
Experiment
In this paper, we utilized the STE Environment (STEenv) as the experimental platform to investigate the application of AGDC in RL. The STE Environment, described in Section 2 or (Shi et al. 2024), comprises a Gaussian diffusion model and a sensor model. Instead of providing rewards to the Agent, it transmits the current position and the test concentration at that position as observations. Any RL algorithm can be integrated with the AGDC module to address the challenges posed by the STE.
Baseline Algorithms and Evaluation Metrics
Infotaxis (Vergassola, Villermaux, and Shraiman 2007) aims to minimize the predicted posterior variance of the source location. It treats the search process as an information-gathering problem, where the objective is to reduce uncertainty about the source location.
Entrotaxis (Hutchinson, Oh, and Chen 2018) enforces the maximum entropy sampling principle, which directs the agent to move towards positions of maximum uncertainty to gather more informative data about the environment and the source location.
DCEE (Chen, Rhodes, and Liu 2021) integrates both exploitation and exploration by incorporating the uncertainty into the control decisions, thus allowing the robot to not only move towards the estimated target but also probe the environment to reduce uncertainty.
@@
Table 1: Parameter Distributions for the Training Scenarios
Scenario Parameterization and Evaluation
@@
Particle Numbers and Cessation Thresholds
@@ Conclusion
Appendix
Experimental details
The policy network structure in our experiments consists of a three-layer fully connected neural network, with each layer containing 128 neurons. This architecture is designed to process complex state inputs, effectively extracting relevant features at each layer to inform action selection. Each layer plays a specific role in refining the input information: the first layer captures essential features from raw state data, the second layer further abstracts these features, and the third layer optimizes the representation for precise action decision-making. This structure enables the network to handle a diverse range of reinforcement learning tasks, facilitating the learning of optimal policies by mapping states to action probabilities efficiently. Additional Hyperparameters in experiments are provided as follows:
Hyperparameter Value Description
$lr$ 1e-4 The learning rate (lr) of DQN
$lr_{actor}$ 1e-4 The lr of actor of DDPG
$lr_{actor}$ 1e-4 The lr of actor of PPO
$lr_{critic}$ 1e-5 The lr of critic of DDPG
$lr_{critic}$ 1e-5 The lr of critic of PPO
Target_update 100 The interval to update network
Lambda ($\lambda$) 0.95 Balancing variance and bias
Minibatch 64 The size of train samples
Replay buffer 1000 The size of replay buffer
discount factor 0.99 The discount factor ($\gamma$)
Table 2: Hyperparameter of RL
Related Work
[…]
Task progress assessment employs various techniques to manage and track complex tasks. Knowledge graphs (Lassila, Hendler, and Berners-Lee 2001; Bollacker et al. 2008) help identify dependencies and bottlenecks, while Bayesian networks (Pearl 1985, 2014) provide probabilistic modeling for assessing task success and risks. Fuzzy control (Zadeh 1973), combined with RL, offers flexible progress estimation in uncertain environments. Hierarchical RL (Parr and Russell 1997; Kulkarni et al. 2016) breaks tasks into sub-goals for easier tracking, and goal management dynamically adjusts priorities to align with overarching objectives. Risk-sensitive RL balances progress with risk for high-stakes decisions, and multi-agent task coordination (Xu et al. 2023b, a, 2022) enables distributed agents to collaboratively advance sub-tasks. Finally, model-based assessment anticipates delays and obstacles, supporting proactive adjustments to ensure efficient task progression.
Normalization Constant in Formula 4
If we do not omit the normalization constant, the weight update formula will include an explicit normalization step. The complete formula is as follows:
1. Compute the unnormalized weight for each particle:
$\bar{w}_{k+1}^{i}=w_{k}^{i}\cdot Pr(o_{k+1}|\Theta_{k+1}^{i})$
2. Compute the sum of the unnormalized weights for all particles:
$C=\sum_{i=1}^{N}\bar{w}_{k+1}^{i}=\sum_{i=1}^{N}\left(w_{k}^{i}\cdot Pr(o_{k+1}|\Theta_{k+1}^{i})\right)$
3. Normalize the weight for each particle:
$w_{k+1}^{i}=\frac{\bar{w}_{k+1}^{i}}{C}=\frac{w_{k}^{i}\cdot Pr(o_{k+1}|\Theta_{k+1}^{i})}{\sum_{j=1}^{N}\left(w_{k}^{j}\cdot Pr(o_{k+1}|\Theta_{k+1}^{j})\right)}$
Therefore, the complete weight update formula (without omitting the normalization constant) is:
$w_{k+1}^{i}=\frac{w_{k}^{i}\cdot Pr(o_{k+1}|\Theta_{k+1}^{i})}{\sum_{j=1}^{N}w_{k}^{j}\cdot Pr(o_{k+1}|\Theta_{k+1}^{j})}$
This form clearly shows how each particle’s weight is normalized to ensure that the sum of all particle weights equals 1.
Baseline algorithms
The notations used in these three algorithms (Infotaxis, Entrotaxis, and DCEE) are summarized as follows: $p_{k}$ represents the robot’s position at time step $k$, $u_{k}$ represents the control input at time step $k$, $\hat{s}_{k}$ represents the estimated source location at time step $k$, $\Theta$ represents the vector of unknown parameters, $Z_{k}$ represents the set of measurements collected up to time step $k$, $b(s)$ represents the belief distribution about the source location given measurements $Z_{k}$, $z_{k}$ represents the measurement taken at time step $k$, $\text{Var}(\hat{s}_{k+1|k})$ represents the variance of the predicted estimation of the source location at time $k+1$, $H(b(s))$ represents the entropy of the belief distribution $b(s)$, $J(u_{k})$ represents the cost function for the respective method, $f(p_{k},u_{k})$ represents the state transition function, predicting the next state based on the current state and control input, $C_{k+1|k}$ represents the predicted covariance matrix of the estimation at time $k+1$, and $\text{trace}(C_{k+1|k})$ represents the trace of the predicted covariance matrix $C_{k+1|k}$.
Infotaxis Infotaxis is an autonomous search strategy where the robot minimizes the predicted variance of the source location estimation. It uses Bayesian inference to update the belief distribution about the source location based on new measurements, continuously refining its understanding of the environment. The core idea is to drive the robot towards actions that maximize the expected information gain, thereby reducing the uncertainty in the source location estimate. This method is particularly useful in environments where the source cannot be directly observed, requiring the robot to infer its location from indirect measurements. Infotaxis has been applied in scenarios such as locating chemical leaks, gas plumes, and even biological searches like finding food sources. This method was proposed by (Vergassola, Villermaux, and Shraiman 2007). in their 2007 paper titled “‘Infotaxis’ as a strategy for searching without gradients”.
Algorithm 2 Infotaxis Algorithm
1:  Initialize robot position $p_{k}$
2:  Initialize belief distribution $b(s)$
3:  while source not located do
4:     Take measurements $z_{k}$
5:     Update belief $b(s)=p(\Theta|Z_{k})$ using Bayesian inference
6:     Define cost function:
7:     $J(u_{k})=\min\text{Var}(\hat{s}_{k+1|k})$
8:     Select control input $u_{k}$ that minimizes the cost function:
9:     $u_{k}=\arg\min_{u}J(u_{k})$
10:     Move robot to new position $p_{k+1}=p_{k}+u_{k}$
11:  end while
Entrotaxis
Algorithm 3 Entrotaxis Algorithm
1:  Initialize robot position $p_{k}$
2:  Initialize belief distribution $b(s)$
3:  while source not located do
4:     Take measurements $z_{k}$
5:     Update belief $b(s)=p(\Theta|Z_{k})$ using Bayesian inference
6:     Define entropy of belief distribution:
7:     $H(b(s))=-\sum_{i}b(s_{i})\log b(s_{i})$
8:     Define cost function:
9:     $J(u_{k})=\max H(b(s))$
10:     Select control input $u_{k}$ that maximizes the cost function:
11:     $u_{k}=\arg\max_{u}J(u_{k})$
12:     Move robot to new position $p_{k+1}=p_{k}+u_{k}$
13:  end while
Entrotaxis is an information-theoretic search method that directs the robot to maximize the entropy of the belief distribution. By focusing on maximizing the information gain, the robot is guided to areas with the highest uncertainty in the current belief. This approach leverages the principle that by exploring these high-uncertainty areas, the robot can collect the most informative data, which significantly enhances the accuracy of the source location estimation. Entrotaxis is especially effective in complex and dynamic environments where the source of the signal is obscured or the signal itself is weak and intermittent. Applications include environmental monitoring, search and rescue operations, and hazardous material detection. This method was detailed by (Hutchinson, Oh, and Chen 2018). in their 2018 paper ”Entrotaxis as a strategy for autonomous search and source reconstruction in turbulent conditions”.
Dual Control for Exploitation and Exploration (DCEE)
Algorithm 4 DCEE Algorithm
1:  Initialize robot position $p_{k}$
2:  Initialize belief about source location $\hat{s}_{k}$
3:  Initialize particle filter with particles $\{\Theta_{i},w_{i}\}$
4:  while source not located do
5:     Take measurements $z_{k}$
6:     Update belief $b(s)=p(\Theta|Z_{k})$ using particle filtering:
7:     $\{\Theta_{i},w_{i}\}\leftarrow\text{ParticleFilterUpdate}(\{\Theta_{i},w_{i}\},z_{k})$
8:     Predict next state using state transition function:
9:     $p_{k+1|k}=f(p_{k},u_{k})$
10:     Update predicted covariance matrix based on current belief and control input:
11:     $C_{k+1|k}=g(b(s),u_{k})$
12:     Define cost function:
13:     $J(u_{k})=\|p_{k+1|k}-\hat{s}_{k+1|k}\|^{2}+\text{trace}(C_{k+1|k})$
14:     Select control input $u_{k}$ that minimizes the cost function:
15:     $u_{k}=\arg\min_{u}J(u_{k})$
16:     Move robot to new position $p_{k+1}=p_{k}+u_{k}$
17:  end while
Dual Control for Exploitation and Exploration (DCEE) is a sophisticated control strategy that balances the dual objectives of exploitation (making use of current best estimates) and exploration (actively reducing uncertainty). DCEE employs a particle filter to update the belief distribution about the source location, incorporating measurements taken by the robot. The cost function in DCEE includes both the distance to the estimated source and the uncertainty in this estimate, quantified by the covariance matrix. This dual consideration ensures that the robot’s actions not only move it towards the likely location of the source but also strategically probe the environment to gather more data, thus refining future estimates. This balanced approach makes DCEE particularly robust and efficient in scenarios where both the environment and source characteristics are uncertain. Applications of DCEE include robotics for environmental monitoring, industrial inspection, and autonomous exploration in unknown terrains. The DCEE method was discussed in the paper (Chen, Rhodes, and Liu 2021) ”Dual Control for Exploitation and Exploration (DCEE) in autonomous search”.
Detailed discussion of the experiment
Analysis Across Various Particle Numbers: The success rates for AGDC-enhanced RL algorithms, including AGDC-DQN, AGDC-PPO, and AGDC-DDPG, show a steady increase as the number of particles increases from 100 to 4000. For instance, AGDC-PPO’s success rate improves from 88.21% at 100 particles to 95.27% at 4000 particles, demonstrating that higher particle numbers enhance the algorithm’s ability to accurately estimate the source by better representing the underlying belief distribution. In contrast, baseline methods like Infotaxis, Entrotaxis, and DCEE also show improved success rates with increased particle numbers, but they consistently lag behind the AGDC-enhanced methods. For example, Infotaxis achieves 88.39% at 4000 particles, which is significantly lower than AGDC-PPO at the same particle count. The Random method remains ineffective, with success rates consistently below 5%, as its performance is not influenced by the number of particles. Note: Unlike the other methods that operate in discrete action spaces, DDPG works in a continuous action space, allowing it to move freely within the environment rather than being restricted to discrete steps along predefined edges.
When analyzing the mean traveled distance, AGDC-enhanced RL algorithms demonstrate a consistent reduction as the number of particles increases. For example, AGDC-DDPG reduces its mean traveled distance from 34.33 at 100 particles to 16.83 at 4000 particles, indicating that more particles allow the algorithms to make more informed decisions, leading to shorter and more efficient paths to the source. In contrast, baseline methods, particularly Entrotaxis, exhibit much higher mean traveled distances, even at higher particle numbers. Entrotaxis starts at 148.83 units with 100 particles and only reduces to 78.69 units at 4000 particles, reflecting the inefficiency of these methods in converging to the source, especially with a lower particle count. The Random method, unaffected by particle count, consistently exhibits very high traveled distances, indicating its lack of efficiency.
In terms of search time, AGDC-enhanced RL algorithms consistently outperform the baseline methods by maintaining significantly shorter search times. AGDC-DQN, for instance, keeps a search time of around 0.94 units at 4000 particles, indicating quick convergence to the source. This efficiency can be attributed to the AGDC module’s ability to guide the search process more effectively as particle numbers increase. On the other hand, baseline methods such as Infotaxis and Entrotaxis show much longer search times, which continue to increase with higher particle numbers. For instance, Infotaxis requires 3.15 units of time at 4000 particles, indicating slower convergence and less effective search strategies compared to the AGDC-enhanced approaches. The Random method, as expected, has minimal search time but fails to produce meaningful results, further underscoring its ineffectiveness.
Analysis Across Various Cessation Thresholds: The success rates of AGDC-enhanced RL algorithms, including AGDC-DQN, AGDC-PPO, and AGDC-DDPG, improve as the cessation thresholds increase, with AGDC-PPO reaching up to 99.13% at a threshold of 0.9. This increase in success rate is due to the algorithms accumulating more confidence before stopping, which results in higher accuracy in locating the source. Baseline methods like Infotaxis and Entrotaxis also show improved success rates with higher cessation thresholds, but their performance remains below that of the AGDC-enhanced methods. Although Infotaxis and Entrotaxis can reach around 99% success rate at the highest thresholds, they require significantly more time and distance to achieve this level of success.
As cessation thresholds increase, AGDC-enhanced RL algorithms demonstrate a decrease in mean traveled distance, indicating more efficient search cessation. For example, AGDC-DDPG reduces its traveled distance to 12.8 units at a threshold of 0.9, reflecting that higher thresholds allow the algorithms to avoid unnecessary exploration and cease the search more efficiently. In contrast, while baseline methods also reduce their traveled distances with higher thresholds, they still require much longer paths compared to AGDC methods. For instance, Infotaxis reduces its traveled distance to 56.4 units at a threshold of 0.9, which is still considerably higher than that of AGDC-DDPG.
Search times for AGDC-enhanced RL algorithms decrease as cessation thresholds increase, with AGDC-PPO reducing its search time to 0.17 units at a threshold of 0.9. This suggests that these algorithms can quickly conclude the search once they have sufficient confidence in their source estimation. However, it is important to note that higher cessation thresholds are not always better. When the threshold exceeds 0.6, the success rate may surpass what would be expected at the upper limit of particle numbers, leading to a misleading improvement in both success rate and distance traveled. This occurs because higher thresholds effectively lower the criteria for task recognition, resulting in the illusion of enhanced performance when in reality, the conditions have simply been made less rigorous.
Descriptions of DQN, PPO, and DDPG
Deep Q-Network (DQN)
Overview: The Deep Q-Network (DQN) is a reinforcement learning algorithm that combines Q-learning with deep neural networks to approximate the Q-value function. It is designed to handle complex decision-making problems in environments with high-dimensional state spaces.
Loss Function: The objective of DQN is to minimize the Bellman error of the Q-value function. The loss function is defined as:
$L(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta^{-})-Q(s,a;\theta)\right)^{2}\right]$
where
$\cdot$ $s$ and $s^{\prime}$ are the current and next states, respectively.
$\cdot$ $a$ and $a^{\prime}$ are the current and next actions, respectively.
$\cdot$ $r$ is the immediate reward.
$\cdot$ $\gamma$ is the discount factor that controls the importance of future rewards.
$\cdot$ $\theta$ represents the parameters of the Q-network.
$\cdot$ $\theta^{-}$ represents the parameters of the target network, which is a delayed copy of the Q-network parameters.
$\cdot$ $\mathcal{D}$ is the experience replay buffer.
Key Features:
Experience Replay: DQN uses a replay buffer to store past experiences and samples batches randomly during training, which breaks the correlation between samples and improves stability.
Target Network: The target network is updated less frequently than the Q-network, reducing oscillations and improving training stability.
$\epsilon$-Greedy Policy: DQN uses an $\epsilon$-greedy policy that gradually reduces the probability of random exploration, thereby improving decision quality over time.
Proximal Policy Optimization (PPO)
Overview: Proximal Policy Optimization (PPO) is an advanced policy gradient method designed to update policies in a stable and efficient manner. PPO achieves this by constraining the magnitude of policy updates, which prevents drastic changes in the policy.
Loss Function: PPO optimizes the policy by minimizing the following clipped objective function:
$L^{\text{CLIP}}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right]$
where
$\cdot$ $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}$ is the ratio between the new policy and the old policy.
$\cdot$ $\hat{A}_{t}$ is the advantage estimate, which measures the relative benefit of taking action $a_{t}$ in state $s_{t}$.
$\cdot$ $\epsilon$ is a hyperparameter that controls the extent of policy updates.
Key Features: $\cdot$ Clipped Surrogate Objective: PPO uses a clipping mechanism to limit the policy update, ensuring that the policy does not change too drastically in a single update, which enhances stability.
$\cdot$ Trust Region Optimization: PPO effectively keeps the new policy close to the old policy, avoiding instability in the policy update process.
$\cdot$ Mini-batch Updates: PPO typically updates the policy using small batches of data multiple times, which improves data efficiency and convergence speed.
Deep Deterministic Policy Gradient (DDPG)
Overview: The Deep Deterministic Policy Gradient (DDPG) is a reinforcement learning algorithm designed for continuous action spaces. It combines the advantages of policy gradient methods and Q-learning to achieve efficient decision-making and control in complex environments.
Loss Functions: DDPG uses two main loss functions for the Critic and Actor networks:
Critic Loss:
$L(\theta^{Q})=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(r+\gamma Q^{\prime}-Q\right)^{2}\right]$
where $Q(s,a;\theta^{Q})$ is the value estimate from the Critic network for the state-action pair $(s,a)$, $Q^{\prime}(s^{\prime},\mu^{\prime}(s^{\prime};\theta^{\mu^{\prime}}(Q^{\prime}));\theta^{Q^{\prime}})$ is the target value estimate from the target Critic network, $Q\mu^{\prime}(s^{\prime};\theta^{\mu^{\prime}})(Q)$ is the action estimated by the target Actor network.
Actor Loss:
$L(\theta^{\mu})=-\mathbb{E}_{s\sim\mathcal{D}}\left[Q(s,\mu(s;\theta^{\mu});\theta^{Q})\right]$
where $\mu(s;\theta^{\mu})$ is the action selected by the Actor network.